OpenStudy (anonymous):

Alfred draws candles randomly from a pack containing 4 colored candles of the same shape and size. There are 2 red candles, 1 green candle, and 1 blue candle. He draws 1 candle and then draws another candle without replacing the first one. Find the probability of picking 1 red candle followed by another red candle, and show the equation used.

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here the probability to draw a red candle as the first candle is: 2/4 = 1/2

OpenStudy (anonymous):

ya i got tht part but the equation part i dont get ^.^

OpenStudy (michele_laino):

now I have to evaluate the probability to draw a red candle as second candle without replacing the first one

OpenStudy (michele_laino):

that probability is 1/3, since after I removed the first red candle there is only one red candle in our pack, and in total there are three candles in our pack

OpenStudy (anonymous):

so what wuld b the equation :s

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

okay thx ^.^

OpenStudy (michele_laino):

I think that the requested probability is given by the multiplication of those probability, namely 1/2 * 1/3 = 1/6

OpenStudy (anonymous):

thank u so much :)

OpenStudy (michele_laino):

thank you! :)